OpenStudy (anonymous):

Evaluate 6 to the power of -3

OpenStudy (anonymous):

evaluate \[\frac{1}{6^3}\]

OpenStudy (anonymous):

no 6^-3

OpenStudy (anonymous):

They are equivalent.

OpenStudy (anonymous):

\[a^{-k}=\frac{1}{a^k}\]

OpenStudy (anonymous):

and the answer is?

OpenStudy (phi):

\[6^{-3}= \frac{1}{6^{3}}=\frac{1}{6\cdot6\cdot6}\] and the answer is?